OpenStudy (anonymous):

12/18 = ?/27 , find the missing number in this proportion

OpenStudy (anonymous):

helpppp

OpenStudy (tkhunny):

Have you considered multiplying by 27?

OpenStudy (tkhunny):

Actually, \(\dfrac{12}{18} = \dfrac{6}{9} = \dfrac{18}{27}\)